# LIO-GVM: an Accurate, Tightly-Coupled Lidar-Inertial Odometry with Gaussian Voxel Map

### v1

@@ IV-C Iterated State Estimation
Upon completion of the IMU and LiDAR process, the subsequent steps for the IESKF-based LIO system involve correspondence matching, residual formulation, and Kalman update. However, the majority of existing filter-based LIO systems [9, 8, 6] typically adopt a straightforward strategy for correspondence matching, that is, matching the point ${}^{L_{k}}\mathbf{p}_{j}$ in the input LiDAR scan with its nearest feature within the target map and treating the ensuing point-to-feature distance as the residual. This strategy involves projecting ${}^{L_{k}}\mathbf{p}_{j}$ to the global frame as:
@@
For long-term odometry, the inevitable accumulation of localization drifts introduces errors to ${}^{G}\bar{\mathbf{T}}_{I_{k-1}}$; Simultaneously, the presence of spike noise, a common issue in less expensive IMUs, leads to inaccuracies in ${}^{I_{k-1}}\hat{\mathbf{T}}_{I_{k}}$. These factors prevent ${}^{L_{k}}\mathbf{p}_{j}$ from being projected to its exact corresponding feature, generating mismatches. Besides, matching the nearest feature in the map could be computationally intensive if the map size grows too large in long-term odometry. To cope with these issues, we propose a new residual formulation for IESKF, which can dismiss incorrectly matched correspondences while maintaining computational efficiency. The details are described in the following subsections.
@@ IV-C1 Correspondence Matching
Figure 3: Neighbor set scheme in implementation given the query voxel ${}^{G}\mathbf{v}_{j}^{n}$.
Suppose the iterated state estimation module is under the $n_{th}$ iteration; the corresponding state estimation will be $\hat{\mathbf{x}}^{n}_{k},\delta\mathbf{x}^{n}_{k}$. We first project the fitted distributions into the global frame as:
@@
${}^{G}\hat{\mathbf{C}}_{j}^{n}$ $={}^{G}\hat{\mathbf{R}}_{L_{k}}^{n}{}^{L_{k}}\mathbf{C}_{j}{}^{G}\hat{\mathbf{T}}_{L_{k}}^{T}.$
For each projected distribution $\left({}^{G}\hat{\boldsymbol{\mu}}_{j}^{n},{}^{G}\hat{\mathbf{C}}_{j}^{n}\right)$, we aim to find its near and similar correspondence in the global map. The global map is composed of voxels stored as a hash table: $\mathbb{Z}^{3}\rightarrow\mathbb{R}^{3}\times\mathbb{R}^{3\times 3}$. The hash table maps the voxel coordinate ${}^{G}\mathbf{v}_{i}$ to its centroid ${}^{G}\boldsymbol{\mu}_{i}$ and the covariance matrix ${}^{G}\mathbf{C}_{i}$ (details of the global map is presented in Section IV-D). In order to evaluate the variance disparity, we define a scalar term derived from the squared Hellinger Distance, which we refer to as the similarity metric $s_{j,i}$:
@@
The larger the $s_{j,i}$ is, the more similar the two distributions are. The correspondence matching procedure can now be conceptualized as follows: We first find the voxel coordinate ${}^{G}\mathbf{v}_{j}^{n}$ in the global map in which ${}^{G}\hat{\boldsymbol{\mu}}_{j}^{n}$ occupies. Then, we define a neighbor set of the occupied voxel as $\mathcal{V}_{j}$ (the selection for $\mathcal{V}_{j}$ in implementation is illustrated in Fig. 3. After that, we calculate the similarity $s_{j,i}$ for all the voxels in $\mathcal{V}_{j}$ and discard the voxel pairs $\left({}^{G}\mathbf{v}_{j}^{n},{}^{G}\mathbf{v}_{i}\right)$ with the similarity smaller than a threshold $s_{t}$. The rest voxel pairs, deemed as effective correspondences, will be utilized for the residual formulation.
@@ IV-C2 Residual Formulation
Before presenting the new residual metric, it is necessary to explain the observation model and state estimate scheme for IESKF. This will provide the rationale behind the proposed residual metric. The observation function can be expressed as:
$\mathbf{z}_{k}^{n}$ $=\mathbf{h}\left({\mathbf{x}}_{k}\right)+\mathbf{v}_{k}$ (12)
$=\mathbf{h}\left(\hat{\mathbf{x}}_{k}^{n}\boxplus\delta\mathbf{x}_{k}^{n}\right)+\mathbf{v}_{k},$
@@
$\mathbf{H}^{n}$ $=\left.\frac{\partial\mathbf{h}\left(\hat{\mathbf{x}}_{k}^{n}\boxplus\delta\mathbf{x}_{k}^{n}\right)}{\partial\delta\mathbf{x}_{k}^{n}}\right|_{\delta\mathbf{x}_{k}^{n}=\mathbf{0}},$ (14)
@@
where $\mathbf{J}^{n}$ is the inverse of the jacobian matrix of $\left(\hat{\mathbf{x}}_{k}^{n}\boxplus\delta\mathbf{x}_{k}^{n}\right)\boxminus\hat{\mathbf{x}}_{k}$ w.r.t. $\mathbf{x}_{k}^{n}$:
$\mathbf{J}^{n}=\left(\left.\frac{\partial\left(\hat{\mathbf{x}}_{k}^{n}\boxplus\delta\mathbf{x}_{k}^{n}\right)\boxminus\hat{\mathbf{x}}_{k}}{\partial\mathbf{x}_{k}^{n}}\right|_{\delta\mathbf{x}_{k}^{n}=\mathbf{0}}\right)^{-1}.$ (16)
Readers may refer to Appendix. -A for the detailed formulation of $\mathbf{J}^{n}$. Combining (13) and (15), following the Bayesian rule, we have the maximize a posterior (MAP) problem:
@@
Given that the matched correspondences are Gaussian distribution pairs, we aim to design a residual metric such that the second term in (17) can represent both the distance and the variance disparity. Let’s consider the weighted Mahalanobis distance presented as follow:
$\sum_{j=0}^{m}(s_{j,i}^{n})^{2}\left({}^{G}\hat{\boldsymbol{\mu}}_{j}^{n}-{}^{G}\boldsymbol{\mu}_{i}\right)^{T}(\hat{\mathbf{C}}_{j}^{n}+\mathbf{C}_{i})^{-1}\left({}^{G}\hat{\boldsymbol{\mu}}_{j}^{n}-{}^{G}\boldsymbol{\mu}_{i}\right),$ (18)
where $m$ is the number of matched correspondences. The Mahalanobis distance is weighted by the similarity of the matched correspondences, ensuring that those with higher similarity and closer distance contribute more to the estimation process, which aligns with our objective. Thus, we recover the second term of (17) back to the non-linear formulation, then equate it to (18). The resulting residual metric for $\left({}^{G}\hat{\boldsymbol{\mu}}_{j}^{n},{}^{G}\hat{\mathbf{C}}_{j}^{n}\right)$ would be:
@@
Please refer to Appendix. -B for the derivation and detailed formulation.
@@ IV-C3 State Update
Please refer to Appendix -C for the derivation details. Suppose we have $m$ correspondences in total, the complete Jacobian matrix is $\mathbf{H}_{k}^{n}=\begin{bmatrix}\mathbf{H}_{k,0}^{n}&\mathbf{H}_{k,1}^{n},...,\mathbf{H}_{k,m}^{n}\end{bmatrix}$. In our implementation, we assume that the covariance matrix $\mathbf{V}_{k}$ of the measurement noise is constant as $\mathbf{V}$. The optimal solution to the MAP problem (17) can be solved by the Kalman update [26]:
$\mathbf{P}^{n}$ $=\mathbf{J}^{n}\hat{\mathbf{P}}_{k}(\mathbf{J}^{n})^{T}$ (21)
$\mathbf{K}^{n}$ $=\left((\mathbf{H}^{n})^{T}\mathbf{V}_{k}^{-1}\mathbf{H}^{n}+(\mathbf{P}^{n})^{-1}\right)^{-1}(\mathbf{H}^{n})^{T}\mathbf{V}_{k}^{-1}$
$\delta\mathbf{x}_{k}^{n+1}$ $=\mathbf{K}^{n}\left(\mathbf{z}_{k}^{n}-\mathbf{h}\left(\hat{\mathbf{x}}^{n}_{k}\right)\right)+\left(\mathbf{K}^{n}\mathbf{H}^{n}-\mathbf{I}\right)\mathbf{J}^{n}(\hat{\mathbf{x}}^{n}_{k}\boxminus\hat{\mathbf{x}}_{k})$
$\hat{\mathbf{x}}_{k}^{n+1}$ $=\hat{\mathbf{x}}_{k}^{n}\boxplus\delta\mathbf{x}_{k}^{n+1}.$
The iteration will stop if the error state $\delta\mathbf{x}^{n+1}_{k}$ satisfies predetermined criteria or the count of iterations surpasses a specified threshold.
Assuming the iteration stops at $n$, we inject $\delta\mathbf{x}_{k}^{n+1}$ into $\hat{\mathbf{x}}_{k}^{n}$ to obtain the estimate of the true state $\bar{\mathbf{x}}_{k}$:
$\hat{\mathbf{x}}_{k}^{n+1}$ $=\hat{\mathbf{x}}_{k}^{n}\boxplus\delta\mathbf{x}_{k}^{n+1},$ (22)
$\bar{\mathbf{x}}_{k}$ $=\hat{\mathbf{x}}_{k}^{n+1}.$
For the covariance matrix, we need to project $\mathbf{P}_{k}$ into the tangent space of $\bar{\mathbf{x}}_{k}$ as:
$\mathbf{P}_{k}=\mathbf{G}\mathbf{P}^{\kappa}\mathbf{G}^{T},\$ (23)
Readers may refer to Appendix. -A for the detailed formulation. The error state can be neglected for major cases, resulting in $\mathbf{G}=\left.\mathbf{I}\right|_{\delta\mathbf{x}_{k}=\mathbf{0}}$; thus many implementations may ignore the state reset for simplification [10, 27]. However, for long-time odometry, adopting this procedure should reduce the accumulated drifts.
@@ IV-D Incremental Global Map
Due to the strategic design of the residual formulation, the global map only necessitates the maintenance of the centroid and covariance matrix for each voxel, which is considerably efficient in terms of memory usage compared to maintaining a tree structure storing all points. Consequently, this enables an incremental update of the global map without the necessity to delete voxels from the map.
@@
The global map is composed of voxels stored as a hash table: $\mathbb{Z}^{3}\rightarrow\mathbb{R}^{3}\times\mathbb{R}^{3\times 3}$. The hash table builds a map from the voxel coordinate ${}^{G}\mathbf{v}_{i}$ to its centroid ${}^{G}\boldsymbol{\mu}_{i}$ and the covariance matrix ${}^{G}\mathbf{C}_{i}$. This structure demonstrates significant efficiency in querying, achieving an optimal $\mathcal{O}(1)$ time complexity for correspondence matching.
The initialization of the global map is similar to the operation in [13]. The first frame is regarded as the global frame $G=L_{0}$. Each point is subsequently fitted to a Gaussian distribution as the procedure in IV-B to obtain $\left({}^{G}\hat{\mathbf{p}}_{j},\mathbf{C}_{j}\right)$. Consequently, the scan is voxelized with voxel size $r$. Suppose we have the voxel with coordinate ${}^{G}\mathbf{v}_{i}$ encompassing a set of points $\{{}^{G}\mathbf{p}_{m}\}$, the centroid and covariance matrix will be calculated as:
${}^{G}\boldsymbol{\mu}_{i}$ $=\frac{1}{M}\sum_{m=0}^{M}{}^{G}\mathbf{p}_{m}$ (24)
${}^{G}\mathbf{C}_{i}$ $=\frac{1}{M}\sum_{m=0}^{M}{}^{G}\mathbf{C}_{m}.$
@@
Once the current estimate $\bar{\mathbf{x}}_{k}$ is acquired, the current scan contributes to the global map. Since the projected Gaussian distributions have been obtained during the state estimation phase, we proceed to voxelize the current scan following (25), resulting in a temporary voxel map ${}^{G}\mathbf{v}_{j}\rightarrow(\bar{\boldsymbol{\mu}}_{j},\bar{\mathbf{C}}_{j})$. For each voxel present in this temporary voxel map, we query its coordinate ${}^{G}\mathbf{v}_{j}$ in the global map. If the global map doesn’t contain the voxel, we directly insert the voxel into the global map. This operation is time-efficient given the insert operation for a hash table has a time complexity of $\mathcal{O}(1)$. Otherwise, the global map will be updated as:
${}^{G}\boldsymbol{\mu}_{j}$ $=\frac{M{}^{G}\boldsymbol{\mu}_{j}+N{}^{G}\bar{\boldsymbol{\mu}}_{j}}{M+N}$ (25)
${}^{G}\mathbf{C}_{j}$ $=\frac{M{}^{G}\mathbf{C}_{j}+N{}^{G}\bar{\mathbf{C}}_{j}}{M+N}$
$M$ $=max(M,N),$
where $M$ is the recorded point number of voxel ${}^{G}\mathbf{v}_{j}$ in the global map, $N$ is the corresponding one in the temporary voxel map.
@@ V Experiment
In this section, the proposed system will be evaluated in terms of accuracy, temporal efficiency, and storage efficiency. All the experiments are conducted on a laptop with a 2.3 GHz AMD Ryzen 7 6800H CPU and 16 GB RAM. We totally choose 9 sequences from two representative datasets for the evaluation. The first 5 sequences, denoted as nc_*, are selected from Newer College dataset [28], with an OS1-64 LiDAR at a scan rate of 10Hz and a built-in 6-axis IMU sampled at 100Hz. The dataset is collected via a handheld device, transported at a moderate walking speed for a distance close to 2.2 kilometers around New College, Oxford. The area covered spans approximately 135$\times$225 $m^{2}$. The rest sequences are selected from MCD VIRAL dataset [29], collected by an OS1-128 LiDAR sampled at 10Hz, an Livox Mid-70 LiDAR and a VectorNav VN100 9-axis IMU sampled at 800 Hz. These sensors are mounted on an All-Terrain-Vehicle (ATV) running at high speed (30 km/h) with aggressive motion around the southern part of the Nanyang Technological University (NTU) Campus, covering around 400$\times$200 $m^{2}$ area. In order to evaluate the influence of IMU on the odometry accuracy, we perform experiments on different IMUs for the same sequences from MCD VIRAL dataset: ntu_* denoting OS1-128 LiDAR with VN100 IMU, ntuo_* denoting OS1-128 LiDAR with its built-in IMU. The ground truth poses for both two datasets are generated by registering each LiDAR scan to a highly accurate prior map using an ICP method.

### v3

@@ IV-C Iterated State Estimation
Upon completion of the IMU and LiDAR processes, the following steps for the IESKF-based LIO system involve correspondence matching, residual formulation, and Kalman update. However, existing filter-based LIO systems [3, 6, 5] typically adopt the nearest correspondence matching strategy and treat the ensuing point-to-feature distance as the residual. This strategy involves projecting ${}^{L_{k}}\mathbf{p}_{j}$ to the global frame as:
@@
For long-term odometry, the inevitable accumulation of localization drifts introduces errors to ${}^{G}\bar{\mathbf{T}}_{I_{k-1}}$. Simultaneously, the presence of spike noise, a common issue in low-cost IMUs, leads to inaccuracies in ${}^{I_{k-1}}\hat{\mathbf{T}}_{I_{k}}$. These factors prevent ${}^{L_{k}}\mathbf{p}_{j}$ from being projected to its exact corresponding feature, generating mismatches (see Fig. 3). Besides, matching the nearest feature in the map could be computationally intensive if the map size grows too large in long-term odometry. To cope with these issues, we propose a new residual formulation for IESKF, which can mitigate incorrectly matched correspondences while maintaining computational efficiency (see Fig. 3). The details are described in the following subsections.
Figure 3: Directly registering the source points to the nearest target points would generate mismatches (shown in the left image). Instead, we: (a). Voxelize the target cloud and fit the points within each voxel with a Gaussian distribution. (b). Fit each source point and its nearest $y$ neighbors with a Gaussian distribution and match it with the near & similar target distributions.
@@ IV-C1 Correspondence Matching
Figure 4: The different neighbor set $\mathcal{V}_{j}$ of a given voxel ${}^{G}\mathbf{v}_{j}^{n}$.
Suppose the iterated state estimation module is under the $n_{th}$ iteration and the corresponding state estimation is $\hat{\mathbf{x}}^{n}_{k},\delta\mathbf{x}^{n}_{k}$. We first project the fitted distributions into the global frame as:
@@
${}^{G}\hat{\mathbf{C}}_{j}^{n}$ $={}^{G}\hat{\mathbf{R}}_{L_{k}}^{n}{}^{L_{k}}\mathbf{C}_{j}{{}^{G}\hat{\mathbf{R}}_{L_{k}}^{n}}^{T}.$
For each projected distribution $\left({}^{G}\hat{\boldsymbol{\mu}}_{j}^{n},{}^{G}\hat{\mathbf{C}}_{j}^{n}\right)$, we aim to find its near and similar correspondence in the global voxel map. Each voxel contains one Gaussian distribution $\left({}^{G}\boldsymbol{\mu}_{i},{}^{G}\mathbf{C}_{i}\right)$ (details of the global map is presented in Section IV-D). Hellinger distance is a statistical measure that quantifies the divergence between two probability distributions. Its value is dominated by two terms: a similarity term and a distance term. The similarity term captures the resemblance between two probability distributions, while the distance term quantifies the separation. For the correspondence evaluation, we extract the similarity term from the squared Hellinger distance [24] and denote it as the similarity metric $s_{j,i}$:
@@
The larger the $s_{j,i}$ is, the more similar the two distributions are. The correspondence matching procedure can now be conceptualized as follows: We first find the voxel coordinate ${}^{G}\mathbf{v}_{j}^{n}$ in the global map in which ${}^{G}\hat{\boldsymbol{\mu}}_{j}^{n}$ occupies. Then, we define a neighbor set of the occupied voxel as $\mathcal{V}_{j}$ (the selection for $\mathcal{V}_{j}$ is illustrated in Fig. 4). After that, we calculate the similarity $s_{j,i}$ between the source distribution and the target distribution in ${}^{G}\mathbf{v}_{i}$ (${}^{G}\mathbf{v}_{i}\in\boldsymbol{\mathcal{V}}_{j}$) and discard the pairs with $s_{j,i}$ smaller than a threshold $s_{t}$. Empirically, the selection for $s_{t}$ is influenced by the noise level of the LiDAR and IMU: the larger the noise is, the smaller the $s_{t}$ should be selected. Generally, $s_{t}$ should not be selected in a strict manner, as the noise inherent in the LiDAR sampling and deskewing procedures could potentially have an adverse impact on similarity.
@@ IV-C2 Residual Formulation
Before presenting the new residual metric, it is necessary to explain the observation model and state estimation scheme for IESKF. This will provide the rationale behind the proposed residual metric. The observation function of the system is:
$\mathbf{z}_{k}^{n}$ $=\mathbf{h}\left(\hat{\mathbf{x}}_{k}^{n}\boxplus\delta\mathbf{x}_{k}^{n}\right)+\mathbf{v}_{k},$ (11)
@@
where $\mathbf{J}^{n}$ is the inverse of the Jacobian matrix of $\left(\hat{\mathbf{x}}_{k}^{n}\boxplus\delta\mathbf{x}_{k}^{n}\right)\boxminus\hat{\mathbf{x}}_{k}$ w.r.t. $\mathbf{x}_{k}^{n}$. Readers may refer to Appendix -A for the detailed formulation of $\mathbf{J}^{n}$. Combining (12) and (13), following the Bayesian rule, we have the maximum a posterior (MAP) problem:
@@
Given that the matched correspondences are Gaussian distribution pairs, we aim to design a residual metric such that the second term in (14) can represent both the distance and the variance disparity. Let’s consider the weighted Mahalanobis distance presented as follows:
$\sum_{j=0}^{m}(s_{j,i}^{n})^{2}\left({}^{G}\hat{\boldsymbol{\mu}}_{j}^{n}-{}^{G}\boldsymbol{\mu}_{i}\right)^{T}(\hat{\mathbf{C}}_{j}^{n}+\mathbf{C}_{i}+\alpha\mathbf{I})^{-1}\left({}^{G}\hat{\boldsymbol{\mu}}_{j}^{n}-{}^{G}\boldsymbol{\mu}_{i}\right),$ (15)
where $m$ is the number of matched correspondences, $\alpha\mathbf{I}$ ($\alpha>0$) is a constant diagonal matrix to avoid the singularity issue. The Mahalanobis distance is weighted by the similarity of the matched correspondences, ensuring that the pairs with higher similarity and closer distance contribute more to the estimation process, which aligns with our objective. Adopting this formula, the resulting residual metric for $\left({}^{G}\hat{\boldsymbol{\mu}}_{j}^{n},{}^{G}\hat{\mathbf{C}}_{j}^{n}\right)$ would be:
@@
Appendix -B exhibits the detailed derivation and definition.
@@ IV-C3 State Update
Please refer to Appendix -A for the derivation and definition details. With the Jacobian matrix, the optimal solution of the MAP problem (14) can be solved by the Kalman update [25] described in Appendix -A.
@@ IV-D Incremental Global Map
The global map is composed of voxels stored as a hash table: $\mathbb{Z}^{3}\rightarrow\mathbb{R}^{3}\times\mathbb{R}^{3\times 3}$, which builds a map from the voxel coordinate ${}^{G}\mathbf{v}_{i}$ to its centroid ${}^{G}\boldsymbol{\mu}_{i}$ and the covariance matrix ${}^{G}\mathbf{C}_{i}$. This structure demonstrates significant efficiency in querying, achieving an optimal $\mathcal{O}(1)$ time complexity for correspondence matching. This structure is similar to that of VGICP [19]. In contrast to VGICP, which only supports pair-wise registration and requires repeated voxelization of the source scan, our approach offers a complete solution for insert, update, and delete operations within a global target map. We also address the variable LiDAR sampling rate problem (e.g. the ground voxels are undersampled in the distance and degrade from the ”plane” to ”line” feature) by introducing an update scheme to ensure the continuous and accurate maintenance of the map.
@@
For the initialization, the first scan is regarded as the global frame $G=L_{0}$. Each point in the scan is subsequently fitted to a Gaussian distribution as the procedure in IV-B to obtain $\left({}^{G}\hat{\mathbf{p}}_{j},\mathbf{C}_{j}\right)$. Consequently, the scan is voxelized with voxel size $r$. Suppose we have the voxel with coordinate ${}^{G}\mathbf{v}_{i}$ encompassing a set of points $\{{}^{G}\mathbf{p}_{m}\}$, the centroid and covariance matrix ${}^{G}\boldsymbol{\mu}_{i},{}^{G}\mathbf{C}_{i}$ will be calculated as the average of the distributions within. Once the current estimate $\bar{\mathbf{x}}_{k}$ is acquired, the current scan contributes to the global map. Since the projected Gaussian distributions have been obtained during the state estimation phase, we proceed to voxelize the current scan following (18), resulting in a temporary voxel map ${}^{G}\mathbf{v}_{j}\rightarrow(\bar{\boldsymbol{\mu}}_{j},\bar{\mathbf{C}}_{j})$. For each voxel present in this temporary voxel map, we query its coordinate ${}^{G}\mathbf{v}_{j}$ in the global map. If the global map doesn’t contain the voxel, we directly insert the voxel into the global map. This operation is time-efficient since the insert operation for a hash table has a time complexity of $\mathcal{O}(1)$. Otherwise, the global map will be updated as:
${}^{G}\boldsymbol{\mu}_{j}$ $=\frac{M{}^{G}\boldsymbol{\mu}_{j}+N{}^{G}\bar{\boldsymbol{\mu}}_{j}}{M+N},$ (18)
${}^{G}\mathbf{C}_{j}$ $=\frac{M{}^{G}\mathbf{C}_{j}+N{}^{G}\bar{\mathbf{C}}_{j}}{M+N},$
where $M$ is the recorded point number of voxel ${}^{G}\mathbf{v}_{j}$ in the global map, $N$ is the corresponding one in the temporary voxel map. In this way, the global map only necessitates the maintenance of the centroid and covariance matrix for each voxel, which is considerably efficient in terms of memory usage and time efficiency compared to maintaining a tree structure storing all points [4] or embedding additional structures into the voxels [5, 16]. Specifically, $M$ is updated using the formula $M=max(M,N)$. This is motivated by the goal of maintaining a continuously updated voxel scheme. This approach ensures that $M$ is constrained by the maximum number of points that can be sampled by LiDAR within a voxel. Otherwise, recalling (18), an unbounded $M$ would lead to $M\gg N$ after a few scans, thus preventing the update of voxel parameters. Consequently, this enables an incremental update of the global map without the necessity to delete voxels from the map. Nevertheless, the mapping scheme also provides the option to remove voxels out of the effective range of the LiDAR for operation in exceptionally large environments.
@@ V Experiment
In this section, the proposed system will be evaluated in terms of accuracy, temporal efficiency, and storage efficiency. All the experiments are conducted on a laptop with a 2.3 GHz AMD Ryzen 7 6800H CPU and 16 GB RAM. We chose 9 sequences from two representative datasets for the evaluation. The first 5 sequences, denoted as nc_*, are selected from Newer College dataset [26], with an OS1-64 LiDAR at a scan rate of 10Hz and a built-in 6-axis IMU sampled at 100Hz. The rest sequences are selected from the MCD VIRAL dataset [27], collected by an OS1-128 LiDAR sampled at 10Hz, a Livox Mid-70 LiDAR, and a VectorNav VN100 9-axis IMU sampled at 800 Hz. In order to evaluate the influence of IMU on the odometry accuracy, we perform experiments on different IMUs for the same sequences from MCD VIRAL dataset: ntu_* denoting OS1-128 LiDAR with VN100 IMU, ntuo_* denoting OS1-128 LiDAR with its built-in IMU. The ground truth poses for both datasets are generated by registering each LiDAR scan to a highly accurate prior map using an ICP method.
ntu_02 ntu_04 ntu_10 ntu_13 ntuo_02 ntuo_04 ntuo_10 ntuo_13 nc_01 nc_02 nc_05 nc_06 nc_07
DLO 6.46/0.43 3.65/0.11 7.65/0.51 3.98/0.20 6.46/0.43 3.75/0.11 7.65/0.51 3.98/0.20 4.20/0.51 4.27/0.37 3.44/0.44 1.31/0.62 1.73/0.18
VoxelMap 1.83/0.11 2.98/0.18 8.48/0.57 $\times$ 1.83/0.11 2.98/0.18 8.48/0.57 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
LIO-SAM 7.80/0.52 3.71/0.23 9.32/0.63 3.05/0.16 -a - - - - - - - -
FAST-LIO2 2.09/0.24 2.11/0.13 2.34/0.15 2.89/0.11 5.64/0.80 3.51/0.23 7.32/0.41 6.79/0.50 3.23/0.37 4.15/0.43 3.13/0.26 0.94/0.37 1.16/0.15
LIO-GVM(w/o s) 2.37/0.40 2.35/0.11 3.21/0.14 2.79/0.14 5.51/0.68 4.73/0.39 6.59/0.39 4.58/0.26 3.36/0.31 4.76/0.44 3.46/0.47 1.28/0.43 1.38/0.20
LIO-GVM(plane) 8.24/0.68 6.91/0.41 8.29/0.74 5.77/0.43 9.34/0.90 7.95/0.54 10.2/0.81 8.89/0.73 6.29/0.49 6.30/0.67 4.54/0.72 2.48/0.59 3.45/0.30
LIO-GVM 1.74/0.46 2.26/0.05 3.16/0.11 2.74/0.09 3.23/0.40 2.69/0.14 4.25/0.27 3.60/0.18 2.47/0.26 3.20/0.35 3.27/0.33 0.89/0.33 1.10/0.17
  • Errors are denoted as ATE[%]/ARE[deg/10m] (smaller is better). We only align the position and orientation at the origin for a comprehensive evaluation.
  • Bold value stands for the best; underline is the second best; ’-’ denotes that the algorithm is not applicable; ’$\times$’ indicates that the algorithm diverges.
  • ’-a’: LIO-SAM is not applicable for ntuo_* an nc_* sequences due to the lack of a 9-axis IMU.